### v2

Abstract
Inference in dynamic probabilistic models is a complex task involving expensive operations. In particular, for Hidden Markov Models, the whole state space has to be enumerated for advancing in time. Even states with negligible probabilities are considered, resulting in computational inefficiency and increased noise due to the propagation of unlikely probability mass. We propose to denoise the future and speed up inference by using only the top-$p$ states, i.e., the most probable states with accumulated probability $p$. We show that the error introduced by using only the top-$p$ states is bound by $p$ and the so-called minimal mixing rate of the underlying model. Moreover, in our empirical evaluation, we show that we can expect speedups of at least an order of magnitude, while the error in terms of total variation distance is below $0.09$.
1 Introduction
Inference in dynamic probabilistic graphical models (dynamic PGMs) [11] is a daunting task as it is in general NP-hard: The probability distribution over all states has to be advanced in time. However, certain future states are highly unlikely. Our goal is to reduce the inference time and denoise the future by only considering the events with a high probability mass. Suppose we have a simple weather model and the sun is shining. In the next time step, the conditions partly cloudy and light rain are much more probable than, e.g., heavy thunderstorm. In this paper, we propose to use only the top-$p$ events, i.e., the most probable events with accumulated probability of at least $p$, for inference, where we only iterate about the more likely states. This way, we reduce the runtime while retaining control over the induced error by setting $p$ appropriately. This approach can also be viewed as a generative AI approach, we generate one possible world, for which we have probability and error assertions.
Let us shortly motivate the runtime savings with language models, e.g., Llama 3 [9]. In language models, we have multibillion parameters and more possible events to be generated than words in the English language. Now, take for example the words The weather is, then the probability of the next word being rainy or sunny is much more probable than eating. With our top-$p$ approach, we can thus spare the iteration over lots of unlikely states. Let us dive a bit deeper into what we actually do by looking at a simple example: Going back to our initial weather model, when we have the weather condition sunny, the next ones could be partly cloudy, light rain, and foggy with relatively high probabilities while a direct transition to more extreme conditions like heavy rain or thunderstorm is less likely. For inference, we now skip the enumeration of heavy rain and thunderstorm. While this change may seem marginal for a minimalistic example at first, it is of great use for very large models: First, the savings are present in each time step, accumulating and reducing the overall time. Second, the model is made sparse, helping human understanding. Also, the obtained sparse HMM can be used in edge computing, where a full HMM may be too big to run.
In our approach, less likely transitions are ignored, leading to potential problems regarding minorities. We note here that the problems arising do not originate from our approach, but from the underlying model. Though, our approach amplifies biases present in the model, e.g., not allowing the switch from sunny to thunderstorm. We also highlight that in applications, where our approach can lead to problems regarding minorities, tools from the field of AI should not be used at all or at least with high caution, as they are prone to biases.
Contribution
In this paper, we propose using only the top-$p$ events for inference in Hidden Markov Models (HMMs) as a representative for PGMs. We show that by using the top-$p$ distributions to proceed in time, we (i) increase the ratio of zeroes in the transition probabilities in our empirical evaluation significantly to above $0.9$, (ii) bound the theoretical error in terms of total variation by $\frac{1-p}{\gamma}$, where $\gamma$ is a model-specific parameter, which we back up in our empirical evaluation by showing that (iii) the introduced error in our examples is only $0.09$ for top-$0.9$ in terms of total variance, and (iv) reduce the runtime for inference by an order of magnitude.
Related Work
Rabiner introduces HMMs [16] and proposes to use the Forward-Backward algorithm [2] for inference. Zhang et al. distill a language model into an HMM [20]. HMMs can be extended to Dynamic Bayesian Networks [7], which can model interactions between different state variables. Inference is done by, e.g, the interface algorithm [14]. However, query answering in Bayesian networks [15] is NP-hard [6], giving rise to approximations. Boyen and Koller represent the distribution over the random variables with successors in the next time step as a product of marginals [5], which is extended by Murphy and Weiss to always use a factored representation [13]. Instead of instantiating a DBN for each time step, Gao et al. introduce a sliding window approach [8]. Murphy provides an overview over approximate inference [14]. Though, the focus in literature is on approximate inference in DBNs or approximation for learning HMMs, but not on approximate inference within HMMs, which we propose in this paper. Vithanage et al. [19] effectively use the $k$ most probable transitions and show that this minimizes the Kullback-Leibler divergence if restricting to $k$ transitions. Contrary, we provide a more fine-grained error assertion.
Structure
The remainder of this paper is structured as follows: First, we introduce HMMs. Second, we explain our top-$p$ model and how to use it. Third, we provide a theoretical analysis of the error and bound the error. Fourth, we provide an empirical evaluation.
2 Hidden Markov Models Generalizing Dynamic Models
In this section, we lay the foundation for our top-$p$ modeling. We introduce HMMs as a representative of dynamic PGMs. An HMM consists of two random variables per time step: One for the state and one for the observation, e.g., the weather condition and whether we see a person wearing a raincoat. The probability of an observation depends only on the current state and the probability of the next state depends only on the current state:
Definition 1(Hidden Markov Model [16])
A Hidden Markov Model consists of two series of random variables $(S_{t},O_{t})_{t}$ over a set $S$ of states and $O$ of observations, where $S_{t}$ denotes the state in time step $t$ and $O_{t}$ the respective observation. The probability $P(O_{t}\mid S_{t})$ of an observation depends only on the current state. The probability distribution over the states is given as the temporal behavior $P(S_{t}\mid S_{t-1})$ and a prior $P(S_{0})$. The semantics for a time step $t$ is thus given by
$P(S_{t}\mid S_{t-1})$ $=\sum_{s_{t-1}\in S_{t-1}}P(S_{t}\mid s_{t-1})\cdot P(s_{t-1})\quad\text{and}$ (1)
$P(O_{t})$ $=\sum_{s_{t}\in S_{t}}P(O_{t}\mid s_{t})\cdot P(s_{t}),$ (2)
where $S_{t}$ is the distribution specifying the probabilities $s_{t}$ for each state $s\in S$.
We illustrate the definition of an HMM with our introductory example:
Example 1
In each time step, the state is the weather condition. The observation is the probability of wearing a raincoat. Table 1 gives the state transition probabilities. The initial state distribution is uniform. The probability of wearing a raincoat is the sum over the rainy conditions light rain, heavy rain, and thunderstorm. When it is sunny, the probability of partly cloudy in the next time step is $0.3$. The probability of wearing a raincoat when it is sunny is $0.35$.
Next/current Partly cloudy Light rain Foggy Sunny Heavy rain Thunderstorm
Partly cloudy 0.3 0.2 0.3 0.3 0.1 0.1
Light rain 0.2 0.2 0.2 0.25 0.2 0.2
Foggy 0.1 0.1 0.2 0.15 0.1 0.1
Sunny 0.2 0.1 0.1 0.2 0.1 0.1
Heavy rain 0.1 0.2 0.1 0.06 0.2 0.3
Thunderstorm 0.1 0.2 0.1 0.04 0.3 0.2
Table 1: State transition probabilities giving the probability of transitioning to the next state given the current state.
Now, the important part is that the sum in Equation 1 goes over all possible states, rendering the equation inefficient for large state spaces, like in language models. Our goal is to reduce the number of states that have to be enumerated. While an HMM seems simple, it can be thought of as a generalized dynamic PGM: We have a, potentially complex, state $S_{i}$ emitting observations $O_{i}$. Let us denote the distributions $P(s_{t}),\,s\in S,$ as forward message updating the state probabilities through time. Dynamic Bayesian Networks (DBNs) [7] can be viewed as HMMs by treating the joint probability distribution over the interface [14] as state and all other random variables jointly as observation. Furthermore, GPT-based language models learn the forward message in a highly complex state space. At their core, they can be viewed as HMMs, interpreting, analogous to DBNs, each possible complex internal state of the language model as a state of an HMM with respective transition and observation (or emission) probabilities. Generating sentences from language models is then the same as predicting states and sampling observations from HMMs, which leads to a noisy distribution in the far future without any observations.
Throughout this paper, we use HMMs as a representative for dynamic PGMs. The same ideas apply to Kalman filters [17] for continuos variables. But for illustrative purposes, we show the idea for discrete variables in an HMM setting. In the next section, we define the top-$p$ approach for HMMs to reduce the space that has to be enumerated for the forward message.
3 Inference in Top-$p$ Hidden Markov Models
Our goal is to denoise the future and reduce the runtime for inference by using only the top-$p$ events, effectively reducing the number of enumerations by not accounting for values not included in the top-$p$ set. Moreover, we focus on the highly probable transitions, denoising the lower ones. With HMMs as our probabilistic model, we present our top-$p$ modeling approach: Given an HMM, we alter the probabilities of the next state given the current state. For all possible next states, we only keep those with the highest probability until we accumulated a total probability mass of $p$. We first formalize this for a single random variable and then extend it to HMMs.
Before we give the definition for top-$p$ HMM, we give a helper definition to select the top probabilities:
Definition 2(Top-$p$ Set)
Given a probability distribution $P$ with a set of events $X$, the sorted order of $P$ is a sequence $Z\subseteq X$ with the events of $X$ arranged descending according to $P(X)$. The top-$p$ set of $P$ is then the set $Y\subseteq X$ built from $Z$ by taking events until their probabilities reach $p$.
For events with equal probability, an arbitrary choice is made. The top-$p$ set is the set of events, also called top-$p$ events, whose probability will be kept and scaled later on, while the probability of events not in the top-$p$ set will be set to zero. We illustrate the sorted order and top-$p$ set with our introductory example:
Example 2
Let $P$ denote the probability of the weather condition in the next time step given it is sunny in the current time step. Table 1 defines the probability distribution $P$. The sorted order of $P$ is then: Partly cloudy, light rain, sunny, foggy, heavy rain and thunderstorm. For $p=0.9$, the top-$p$ set consists of partly cloudy, light rain, foggy and sunny. For $p=0.91$, the top-$p$ set additionally includes heavy rain.
Weather condition Probability
Partly cloudy $\frac{1}{3}$
Light rain $\frac{5}{18}$
Foggy $\frac{1}{6}$
Sunny $\frac{2}{9}$
Heavy rain 0
Thunderstorm 0
Table 2: Top-$p$ distribution for the state in the next time step given it is currently sunny for $p=0.9$.
Building on the top-$p$ set, we define the top-$p$ distribution:
Definition 3(Top-$p$ Distribution)
Given a probability distribution $P$ with a set of events $X$, the top-$p$ distribution of $P$ is another probability distribution $Q$ over the same set of events $X$. Let $Y$ be the top-$p$ set of $P$ for a given $p$. Then, $Q(X)$ is defined as
$Q(x)=\begin{cases}\frac{P(x)}{P(Y)}&x\in Y\\ 0&\text{otherwise}\end{cases}\ .$ (3)
In other words, the probabilities for all events not in the top-$p$ probabilities are set to zero while the other probabilities are scaled to sum up to one. For Example 2 and $p=0.9$, Table 2 shows the top-$p$ distribution. As an HMM consists of conditional probability distributions, we independently apply the top-$p$ distributions to them:
Definition 4(Top-$p$ HMM)
Given an HMM $H$ with a set $S$ of states, a set $O$ of observations, probability distributions over the states given by $P_{H}(S_{0})$ and $P_{H}(S_{i+1}\mid s_{i})$, for each state $s\in S$, and probability distributions $P_{H}(O_{i}\mid s_{i})$, $s\in S$, over the observations, the top-$p$ HMM of $H$ is another HMM $Q$ over the same set $S$ of states and same set $O$ of observations. The probability distributions of $Q$ are the top-$p$ distributions of the respective ones of $H$, that is, (i) $P_{Q}(S_{0})$is the top-$p$ distribution of $P_{H}(S_{0})$, (ii) $P_{Q}(S_{i+1}\mid s_{i})$is the top-$p$ distribution of $P_{Q}(S_{i+1}\mid s_{i})$, for each state $s\in S$, and (iii) $P_{Q}(O_{i}\mid s_{i})$is the top-$p$ distribution of $P_{H}(O_{i}\mid s_{i})$, $s_{i}\in S$.
Algorithm 1 is pseudo-code for Definition 3 and returns the top-$p$ distribution for a vector representing a probability distribution. For an HMM, Algorithm 1 is applied to (i) the initial state distribution, (ii) each next state distribution given the current state, (iii) and each observation distribution given the current state. We illustrate the top-$p$ HMM on our running example:
Algorithm 1 Apply the top-$p$ approach according to Definition 3 to a vector
$v\in[0,1]^{d},p\in(0,1]$ $\triangleright$ Input probability distribution and $p$ value
$v^{\prime}\in[0,1]^{d}$ is the top-$p$ distribution of $v$
$v^{\prime}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$
$idx\leftarrow\text{sort\_indices\_by\_value\_desc}(v)$ $\triangleright$ create sorted order of $v$
$sum\leftarrow 0$
for $j$ in idxidx do $\triangleright$ $j$ is an index for $v$ and $v^{\prime}$
  $sum\leftarrow sum+v[j]$
  $v^{\prime}[j]\leftarrow v[j]$ $\triangleright$ take events …
  if $sum\geq p$ then $\triangleright$ …until accumulated probability reaches $p$
   break
  end if
end for
$v^{\prime}\leftarrow\frac{v^{\prime}}{sum}$ $\triangleright$ scale to sum to one
Example 3
Table 3 shows the state transition probabilities for $p=0.7$. For the observations, the probability of wearing a rain coat is one for heavy rain and thunderstorm and unaltered for all other states.
Next/current Partly cloudy Light rain Foggy Sunny Heavy rain Thunderstorm
Partly cloudy $\frac{3}{7}$ 0.25 $\frac{3}{7}$ 0.4 0 0
Light rain $\frac{2}{7}$ 0.25 $\frac{2}{7}$ $\frac{1}{3}$ $\frac{2}{7}$ $\frac{2}{7}$
Foggy 0 0 $\frac{2}{7}$ 0 0 0
Sunny $\frac{2}{7}$ 0 0 $\frac{4}{15}$ 0 0
Heavy rain 0 0.25 0 0 $\frac{2}{7}$ $\frac{3}{7}$
Thunderstorm 0 0.25 0 0 $\frac{3}{7}$ $\frac{2}{7}$
Table 3: Top-$p$ state distributions for the simple weather HMM with $p=0.7$.
Having shifted the probability mass, we now exploit the new distributions for efficient inference without enumerating all states. Since the transition matrix still has the same dimensions, but is more spare, we use sparse matrices to not enumerate all states. We explain how to perform inference, by (sparse) matrix multiplication, in a top-$p$ HMM $Q$ of an original HMM: (1) The top-$p$ HMM $Q$ is built according to Definition 4 by using Algorithm 1. (2) The obtained transition probabilities and observation probabilities are stored in a transition and observation matrix, respectively. (3) The two matrices are each transformed into compressed row storage [1]: For each row in the matrix, a tuple $(c,x)$ for each non-zero entry, where $c$ is the column number and $x$ the value. (4) For inference, sparse matrix-vector multiplication [4] is used to proceed the current state distribution in time and to get the observation distribution: The dot product between each row and the vector is calculated by summing the products of $x$ and the entry $c$ of the vector for each $(c,x)$ tuple in the row. Sparse matrix-vector multiplication runs in time proportional to the number of non-zero entries [4]. Sparse matrix-vector multiplication is also available for GPUs [3].
Since we only keep the top-$p$ events, our top-$p$ HMM definitely introduces an error. For example, in Table 3 we see that we introduce an error by dropping the possibilities of transitioning from sunny to heavy rain or thunderstorm. Next, we show that the error induced by top-$p$ modeling is bounded in terms of $p$.
4 Bounding the Approximation Error
When we use the top-$p$ HMM, inference is done with altered distributions, dropping events of the original one. In this section, we bound the approximation error introduced by the top-p modeling. The bound can then be used to determine whether to apply the top-$p$ approach or not. Throughout this section, we use the total variation distance to measure the approximation error:
Definition 5(Total Variation)
The total variation distance between two discrete probability distributions $P$ and $Q$ over the same set of possible outcomes $X$ is defined as
$\delta(P,Q)=\frac{1}{2}\sum_{x\in X}\left\lvert P(x)-Q(x)\right\rvert$. (4)
We choose the total variation distance, because it allows, in contrast to the Kullback–Leibler divergence [12], arbitrary probabilities being zero. Moreover, the absolute value can be more easily resolved than, e.g., square roots in the Helliner distance and the total variation fulfills the triangle equality. We first start by bounding the error of a single time step approximation and continue to bound the error over all time steps based on the result for one time step.
4.1 Approximation Error in a Single Time Step
We show that the error introduced by a top-$p$ distribution is at most 1p1-p:
Theorem 4.1
Given a probability distribution $P$ and the top-$p$ distribution $Q$ of $P$, the total variation distance between $P$ and $Q$ is bounded by
$\delta(P,Q)\leq 1-p.$ (5)
Proof
Let us denote with $Y$ the top-$p$ set of $P$. We first assume that the probabilities kept exactly match $p$, i.e., $P(Y)=p$ and $Q(Y)=1$. Moreover, for each $y\in Y$: $Q(y)=\frac{P(y)}{p}$. Now, we can start deriving the distance:
$\delta(P,Q)$ $=\frac{1}{2}\sum_{x\not\in Y}\left\lvert P(x)\right\rvert+\frac{1}{2}\sum_{x\in Y}\left\lvert P(x)-Q(x)\right\rvert$ (6)
$=\frac{1}{2}(1-p)+\frac{1}{2}\sum_{x\in Y}\left\lvert P(X)-\frac{P(x)}{p}\right\rvert$ (7)
$=\frac{1}{2}(1-p)+\frac{1}{2}\sum_{x\in Y}\left\lvert P(X)\left(1-\frac{1}{p}\right)\right\rvert$ (8)
$=\frac{1}{2}(1-p)+\frac{1}{2}\sum_{x\in Y}\left\lvert P(x)\right\rvert\left\lvert 1-\frac{1}{p}\right\rvert$ (9)
$=\frac{1}{2}(1-p)-\left(1-\frac{1}{p}\right)\cdot\frac{1}{2}\sum_{x\in Y}\left\lvert P(x)\right\rvert$ (10)
$=\frac{1}{2}(1-p)-\left(1-\frac{1}{p}\right)\cdot\frac{1}{2}p=1-p.$ (11)
Lifting the assumption of $P(Y)=p$, if the probabilities in the top-$p$ set exceed $p$, i.e., $P(Y)=p^{\prime}\geq p$, we have $\delta(P,Q)=1-p^{\prime}\leq 1-p$. ∎
Because the total variation distance is invariant to linear transformations and by the triangle inequality, we can give a first bound on multiple steps:
Corollary 1
Given an HMM with a probability distribution $P^{k}$ over the set of states for each time step $k$ and a probability distribution $Q^{k}$, which is the top-$p$ distribution of the previous $Q^{k-1}$ and $Q^{0}$ is the top-$p$ distribution from $P^{0}$, then the total variation distance is
$\delta(P^{k},Q^{k})\leq(k+1)\cdot(1-p).$ (12)
Unfortunately, the bound in Equation 12 quickly approaches the trivial bound of $1$. Boyen and Koller show that the error in terms of Kullback-Leibler divergence is bounded [5]. However, the Kullback-Leibler divergence is not applicable to our top-$p$ HMM, because we set probabilities to zero and the Kullback-Leibler divergence divides by probabilities. Nevertheless, we use the key idea of Boyen and Koller [5] to show the same bound for our total variance distance.
4.2 Overall Approximation Error
Boyen and Koller show for the Kullback-Leibler divergence that the error in each time step is at most $\frac{\epsilon}{\gamma}$, where $\gamma$ is a parameter inherent to the specific HMM and $\epsilon$ is the error incurred in each time step [5]. The parameter $\gamma$ of an HMM tells the minimum probability mass on which two different state distributions agree on after a single time step. While the bounded error result is promising, we cannot use their theorems directly, because of their use of the Kullback-Leibler divergence, which is not applicable in our case. The Kullback-Leibler divergence divides by probabilities and we set probabilities to zero, rendering the Kullback-Leibler divergence useless in our case. Therefore, we use the key idea from Boyen and Koller [5] to bound the error our top-$p$ HMM in terms of our used total variation distance. We start by defining the parameter $\gamma$ and continue by proving that an error incurred in one time step diminishes by a factor of $1-\gamma$ over the next time steps. Then, given that our error increases at most by 1p1-p in each time step (c.f. Corollary 1), we build a geometric series and prove that our overall approximation error in terms of total variation distance is at most $\frac{1-p}{\gamma}$ in each time step.
Let us first prove that an error introduced in one time step diminishes by a constant factor per time step. For this, we define the minimal mixing rate $\gamma$, which directly links to the diminishing factor of $1-\gamma$. The minimal mixing rate sets the minimum probability mass two different state distributions agree on after a single time step [5].
Definition 6(Definition 3 from Boyen and Koller [5])
For a Markov process with stochastic transition model $Q$ with states $\omega_{i}$, the minimal mixing rate of $Q$ is
$\gamma\coloneq\min_{i_{1},i_{2}}\sum_{j}\min\left\{Q\left(\omega_{j}\mid\omega_{i_{1}}\right),Q\left(\omega_{j}\mid\omega_{i_{2}}\right)\right\}.$ (13)
For showing the diminishing factor of $1-\gamma$ for the error, we split the transition from one time step to the next into two transitions containing an intermediate state. We are primarily interested in the first transition, as we show later, using the following lemma, that this one diminishes the error.
Lemma 1(Lemma 2 and Theorem 3 by Boyen and Koller [5])
Fixing two state distributions $\varphi$ and $\psi$, the transition of a stochastic process $Q$ can be split into two steps: First, a transition defined by $R^{\Gamma}$ from $\Omega$ to $\tilde{\Omega}$ and, second, a transition defined by $R^{\Delta}$ from $\tilde{\Omega}$ to $\Omega$, where $\Omega=\left\{\omega_{i}\right\}_{i}$ is the state space of $Q$ and $\tilde{\Omega}=\Omega\cup\{c\}$ with a new intermediate state $c$. The process $R^{\Gamma}$ preserves its state with probability $1-\gamma$ and transitions to state $c$ with probability $\gamma$. The process $R^{\Delta}$ transitions from $\omega_{i}$ to $\omega_{j}$ with probability $\frac{Q_{i,j}^{\Delta}}{1-\gamma}$. For state $c$, $R^{\Delta}$ it transitions to state $\omega_{j}$ with probability $\sum_{i}\varphi(\omega_{i})\frac{Q_{i,j}^{\Gamma}}{\gamma}$. The matrices $Q^{\Gamma}$ and $Q^{\Delta}$ arise from an additive contraction decomposition $Q=Q^{\Gamma}+Q^{\Delta}$, where $Q^{\Gamma}$ and $Q^{\Delta}$ are non-negative matrices such that, for all $i$, $\sum_{j}Q_{i,j}^{\Gamma}=\gamma$, and for all $j$, $\sum_{i}\varphi(\omega_{i})Q_{i,j}^{\Gamma}=\sum_{i}\psi(\omega_{i})Q_{i,j}^{\Gamma}$.
We now show that the first transition diminishes the error by $1-\gamma$ and that the second one does not increase the error:
Theorem 4.2
For two state distributions $\varphi$ and $\psi$ and their counterparts $\varphi^{\prime}$ and $\psi^{\prime}$ in the next time step as defined by a stochastic process $Q$:
$\delta(\varphi^{\prime},\psi^{\prime})\leq\left(1-\gamma_{Q}\right)\delta(\varphi,\psi).$ (14)
Proof
[…]
For showing $\delta(\tilde{\varphi},\tilde{\psi})=(1-\gamma)\delta(\varphi,\psi)$, we first note
$\tilde{\varphi}(c)$ $=\sum_{i}\gamma\varphi(\omega_{i})=\gamma\quad\text{and}$ (15)
$\tilde{\varphi}(\omega_{i})$ $=(1-\gamma)\varphi(\omega_{i}),$ (16)
which also hold for $\tilde{\psi}$. Then, we have
[…]
$=\frac{1}{2}\sum_{i}\left\lvert\left(1-\gamma\right)\varphi(\omega_{i})-\left(1-\gamma\right)\psi(\omega_{i})\right\rvert$ (18)
$=\left(1-\gamma\right)\delta(\varphi,\psi).$ (19)
At this point, we have already shown that we (i) introduce an approximation error of 1p1-p when using the top-$p$ distribution once (c.f. Theorem 4.1), (ii) the error introduced in one time step diminishes by a factor of $1-\gamma$ for each future time step (c.f. Theorem 4.2), which is $(1-p)(1-\gamma)^{k}$ for time step $k$, and (iii) in each time step we only add an additional error of at most 1p1-p (c.f. Corollary 1). We combine these results in our overall approximation guarantee by analyzing the induced geometric series:
Theorem 4.3
The total variation distance between the state distributions for the top-$p$ HMM $Q$ and the original HMM $P$ in each time step $k$ is
$\delta(P^{k},Q^{k})\leq\frac{1-p}{\gamma},$ (20)
where $\gamma$ is the minimal mixing rate of $P$.
Proof
We prove the claim by induction: In the first time step, the error is $1-p\leq\frac{1-p}{\gamma}$ by Theorem 4.1 and because of $\gamma\leq 1$.
Induction Step
In time step $k$, because of the triangle inequality, the error consists of the error from time step k1k-1 and an additional error of at most 1p1-p. However, the error of time step k1k-1 is, by Theorem 4.2, diminished by the factor $1-\gamma$. Therefore, the error in time step $k$ is
$(1-\gamma)\cdot\frac{1-p}{\gamma}+1-p$ $=(1-\gamma)\cdot\frac{1-p}{\gamma}+\gamma\cdot\frac{1-p}{\gamma}$ (21)
$=(1-\gamma+\gamma)\cdot\frac{1-p}{\gamma}=\frac{1-p}{\gamma}.$ (22)
We highlight that Theorem 4.3 helps an agent to decide whether to use the top-$p$ approach beforehand based on the error bound. Moreover, the exact total variation can be calculated for individual HMMs. In the next section, we substantiate our theoretical margin of error empirically, show a reduced runtime for inference, and an increased sparsity.
5 Evaluation
We evaluate the top-$p$ approach on three HMMs: One favoring our approach, one that is disadvantageous for the top-$p$ approach and a language model for a realistic application. We investigate the sparsity, the runtime with and without observations, and the total variation distance of the top-$p$ model. We have trained a very simple language model as an HMM with 7620 states, which we call LM HMM. In practice, one would distill a larger model into an HMM, like in Zhang et al. [20]. In this evaluation, we focus on the runtime and not on the quality of the language model. We create a synthetic Bell HMM, which is designed to favor the top-$p$ approach. The Bell HMM distributes, in each next-state distribution, a probability mass of 0.9 over five states and the remaining 0.1 over 795 states, favoring our approach by having lots of unlikely successor states. For a synthetic HMM disadvantageous for the top-$p$ approach, we create a Uniform HMM consisting of uniform state distributions for a system of 800 states. The number of observations equals the number of states in all HMMs. We evaluate top-$p$ for the $p$ values $0.5$, $0.7$, and $0.9$ and for the maximum time step $50$. We mostly give only the numbers for $p=0.9$ as the three top-$p$ approaches are mostly comparable. We run all tests on a 13th Gen Intel(R) Core(TM) i5-1345U CPU with 1.60 GHz and 16 GB of RAM. We use NumPy [10] and SciPy [18] for matrix and sparse matrix operations, respectively.
Sparsity We define the sparsity as the ratio of zero-entries in the transition matrix. In the Bell HMM, all three top-$p$ approaches lead to a sparsity of above $0.99$. In the Uniform HMM, the sparsity is approximately 1p1-p. And in the LM HMM, the sparsity is for all three top-$p$ greater than $0.97$.
Runtime for Inference The top-$0.9$ HMM for the Bell HMM outperforms exact inference in the original HMM by a factor of more than $15$. For the Uniform HMM, top-$0.9$ is around $3.39$ times slower than the original HMM. The reason for the slower runtime is the low sparsity: Sparse matrix multiplication only pays off for high sparsity. In the LM HMM, top-$p$ is more than $18$ times faster than the original HMM. Figure 1(a) shows the experiment for the LM HMM. In Appendix 0.A.1, we show the experiment for the Bell and Uniform HMM.
Runtime for Filtering Before, we reported runtimes without observations. Now, we enter an observation every five time steps. We omit the LM HMM, since observation in text generation is not applicable. In the Bell HMM, the top-$0.9$ approach is more than $77$ times faster than filtering in the original HMM. In the Uniform HMM, the top-$0.9$ approach is more than $24$ times faster than the original HMM. Figure 1(b) shows the experiment for the Bell HMM. In Appendix 0.A.1, we show the experiment for the Uniform HMM.
Total Variation Distance The total variation in the Bell HMM is $0.086$ for top-$0.9$, $0.285$ for top-$0.7$ and $0.78$ for top-$0.5$. In the Uniform HMM, the total variation distance is $0.099$, $0.299$ and $0.499$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively. In the LM HMM, the total variation distance is $0.059$, $0.172$ and $0.3$ for top-$0.9$, top-$0.7$ and top-$0.5$, respectively.
Total Variation Distance for Filtering With observations, the total variation in the Bell HMM is $0.095$ for top-$0.9$, $0.491$ for top-$0.7$ and $0.846$ for top-$0.5$. In the Uniform HMM, the total variation distance is the same as for no observations, since the model is uniform. We again omit the LM HMM, since observation in text generation is not applicable.
(a) LM HMM
(b) Bell HMM with filtering
Figure 1: Runtime for inference in the LM and Bell HMMs. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Summing up, the top-$p$ approach significantly increases sparsity to more than $0.9$ for top-$0.9$. Moreover, top-$p$ is at least an order of magnitude faster than the original HMMs, while the total variation distance for top-$0.9$ stays below $0.09$.
6 Conclusion
Inference in large (dynamic) probabilistic models like HMMs, DBNs, or GPT-based language models like Llama 3, is a complex task involving expensive operations. In particular, the whole state space needs to be enumerated for advancing in time. We propose to only use the top-$p$ events, i.e., the most probable events with an accumulated probability of $p$. Using only the top-$p$ events and setting the probability for all other ones to zero, we can significantly speedup inference by at least an order of a magnitude. Moreover, the error introduced by the top-$p$ HMM can be bound in terms of total variation distance. For a simple language model, the total variation distance is below $0.059$ for top-$0.9$. Furthermore, an agent can decide whether to use the top-$p$ approach by checking the error bound. In future work, the top-$p$ approach can be utilized for first-order inference in DBNs. Moreover, the obtained sparse top-$p$ HMM can be used in edge computing, where a full HMM may be too big to run on.
{credits}
6.0.1 Acknowledgements
The research for this paper was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2176 ’Understanding Written Artefacts: Material, Interaction and Transmission in Manuscript Cultures’, project no. 390893796. The research was conducted within the scope of the Centre for the Study of Manuscript Cultures (CSMC) at Universität Hamburg.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Appendix
0.A.1 Evaluation
In this appendix, we show more plots for the evaluation in Section 5.
Runtime for Inference
Figure 2 shows the runtime for inference in the Bell HMMs. All three top-$p$ approaches are significantly faster than inference in the original HMM and are comparable fast. Figure 3 shows the runtime for inference in the Uniform HMM. Clearly, the lower $p$, the faster the inference. However, all top-$p$ models have a slower runtime than the original one. Both is because the sparsity increases with higher $p$ and sparse matrix multiplication only pays off when the sparsity is high.
Figure 2: Runtime for inference in the Bell HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Figure 3: Runtime for inference in the Uniform HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.
Runtime for Filtering
Figure 4 shows the runtime for filtering every five steps. In contrast to regular prediction, the top-$p$ approaches are now significantly faster than filtering in the original HMM.
Figure 4: Runtime for filtering in the Uniform HMM. The blue line shows the runtime in the original HMM and the other three ones for different top-$p$ ones.